What is the pH of a solution that is 0.785 M in acetic acid and 0.451 M is sodium acetate. The Ka for acetic acid is 1.8 x 10-5 .

A.

4.74

B.

4.50

C.

4.74

D.

4.98

E.

5.50
Answer #1

Ka = 1.8 * 10^-5

PKa = -log(Ka)

= -log(1.8 * 10^-5)

= 4.74

Ph = Pka + log (Salt /Acid)

Ph = 4.74 + log(0.451 / 0.785)

= 4.74 + log(0.574)

= 4.74 - 0.24

= 4.50

Correct answer :- B